A 1.00 L solution contains 0.352 M HA (Ka = 4.66⋅10−64.66⋅10-6) and 0.758 M NaA.

What is the pH of this solution?
What is the pH if 0.310 L of 0.138 M KOH is added to the solution be?
What is the pH if 0.620 L of 0.138 M KOH is added to the solution be?
Answer #1

Given solution contain weak acid HA and its salt NaA. Therefore this solution is a buffer solution. We can use Henderson's equation to calculate pH of buffer solution.

We have Henderson's equation, pH = pKa + log [ Salt ] / [ acid ]

\therefore pH = - log Ka + log [ NaA ] / [ HA ]

\therefore pH = - log 4.66 \times 10 -06 + log 0.758 / 0.352

pH = 5.332 + log 0.758 / 0.352

pH = 5.332 + 0.333

pH = 5.665

Question 2

Calculation of moles of NaA, HA and KOH.

We have relation, Molarity = No. of moles of solute / Volume of solution in L

No. of moles of solute = Molarity \times volume of solution in L

No. of moles of HA = 0.352 mol / L \times 1.00 L = 0.352 mol

No. of moles of NaA = 0.758 mol / L \times 1.00 L = 0.758 mol

No. of moles of KOH = 0.138 mol / L \times 0.310 L = 0.04278 mol

Consider reaction of KOH with buffer solution.

HA + OH -  \rightarrow A - + H2O

Let's use ICE table.

Moles HA OH - A -
I 0.352 0.04278 0.758
C -0.04278 -0.04278 +0.04278
E 0.30922 0.00000 0.80078

Volume of buffer solution after addition of KOH = 1.00 L + 0.310 L = 1.310 L

We have equation, pH = - log Ka + log [ NaA ] / [ HA ]

\therefore pH = - log 4.66 \times 10 -06 + log ( 0.80078 mol / 1.310 L ) / ( 0.30922 mol / 1.310 L )

pH = 5.332 + 0.4132

pH =5.745

Question 3

Calculation of moles of NaA, HA and KOH.

We have relation, Molarity = No. of moles of solute / Volume of solution in L

No. of moles of solute = Molarity \times volume of solution in L

No. of moles of HA = 0.352 mol / L \times 1.00 L = 0.352 mol

No. of moles of NaA = 0.758 mol / L \times 1.00 L = 0.758 mol

No. of moles of KOH = 0.138 mol / L \times 0.620 L = 0.08556 mol

Consider reaction of KOH with buffer solution.

HA + OH -  \rightarrow A - + H2O

Let's use ICE table.

Moles HA OH -   A -
I 0.352 0.08556 0.758
C -0.08556 -0.08556 +0.08556
E 0.26644 0.00000 0.84356

Volume of buffer solution after addition of KOH = 1.00 L + 0.620 L = 1.620 L

We have equation, pH = - log Ka + log [ NaA ] / [ HA ]

\therefore pH = - log 4.66 \times 10 -06 + log ( 0.84356 mol / 1.620 L ) / ( 0.26644 mol / 1.620 L )

pH = 5.332 + 0.5005

pH =5.832